for the straight line defined by the point 3, 59 and 5, 87 determine the slope and y intercept

Answer by Boreal(15235)

slope is 28/2=14
point slope formula is y-59=14(x-3)
y-59=14x-42
y=14x+17
14 is the slope, 17 the y-intercept.